OpenStudy (anonymous):

f(x,t)=arctan(xsqurt(t)) dervived to the respect of x and then t. Partial Derivatives

OpenStudy (anonymous):

\[f _{x}(x,t)=\sqrt{t}/(x ^{2}t+1)\]\[f _{xt}(x,t)=((x ^{2}t+1)/(2\sqrt{t})-x ^{2}\sqrt{t}))/(x ^{2}t+1)^{?}\]

OpenStudy (anonymous):

? will be 2

OpenStudy (anonymous):

i need to see some steps. f(x,t) to the respect of t. so i know x is a constant but can you take me farther. I was wanting you to show me how not tell me the answer

OpenStudy (anonymous):

ok.......

OpenStudy (anonymous):

\[df/dx =[1/((x \sqrt{t})^{2}+1)]d(x \sqrt{t})/dx=\sqrt{t}/(x ^{2}t+1)\]\[df _{x}/dt=[(x ^{2}t+1)d(\sqrt{t})/dt-\sqrt{t}d(x ^{2}t+1)/dt)](x ^{2}t+1)^{2}\]

OpenStudy (anonymous):

this gives you the ans.

OpenStudy (anonymous):

thanks